# Resilience of Entropy Model in Distributed Neural Networks

### v1

@@ Abstract
Distributed deep neural networks (DNNs) have emerged as a key technique to reduce communication overhead without sacrificing performance in edge computing systems. Recently, entropy coding has been introduced to further reduce the communication overhead. The key idea is to train the distributed DNN jointly with an entropy model, which is used as side information during inference time to adaptively encode latent representations into bit streams with variable length. To the best of our knowledge, the resilience of entropy models is yet to be investigated. As such, in this paper we formulate and investigate the resilience of entropy models to intentional interference (e.g., adversarial attacks) and unintentional interference (e.g., weather changes and motion blur). Through an extensive experimental campaign with 3 different DNN architectures, 2 entropy models and 4 rate-distortion trade-off factors, we demonstrate that the entropy attacks can increase the communication overhead by up to 95%. By separating compression features in frequency and spatial domain, we propose a new defense mechanism that can reduce the transmission overhead of the attacked input by about 9% compared to unperturbed data, with only about 2% accuracy loss. Importantly, the proposed defense mechanism is a standalone approach which can be applied in conjunction with approaches such as adversarial training to further improve robustness. Code will be shared for reproducibility.
@@ 2 Entropy Coding in Distributed DNNs
Recent work [41] has incorporated the above-mentioned lossless coding with lossy compression technique into distributed DNNs to increase communication efficiency between mobile and edge devices. The proposed method, referred to as the entropic student, aims at achieving a balance between the coding rate and task-specific performance metrics (e.g., cross entropy in classification tasks). One distinction between the entropic student and other entropy-coding-based feature compression work [54, 58] lies in the significantly smaller encoder size of the former due to constraints on computational resources in mobile devices. As a result, the entropic student splits the DNN at an early layer and use knowledge distillation [25] for training the head to preserve accuracy. However, existing work does not address the issue of resilience of entropy models, which is the key goal of this paper.
@@ 3 Modeling Interference to Entropy Coding
$\min_{\delta}\quad$ $\log_{2}P_{Z}(H(x+\delta))$ (7)
$s.t.\quad$ $||\delta||_{p}\leq\epsilon$

### v2

@@ Abstract
Distributed deep neural networks (DNNs) have emerged as a key technique to reduce communication overhead without sacrificing performance in edge computing systems. Recently, entropy coding has been introduced to further reduce the communication overhead. The key idea is to train the distributed DNN jointly with an entropy model, which is used as side information during inference time to adaptively encode latent representations into bit streams with variable length. To the best of our knowledge, the resilience of entropy models is yet to be investigated. As such, in this paper we formulate and investigate the resilience of entropy models to intentional interference (e.g., adversarial attacks) and unintentional interference (e.g., weather changes and motion blur). Through an extensive experimental campaign with 3 different DNN architectures, 2 entropy models and 4 rate-distortion trade-off factors, we demonstrate that the entropy attacks can increase the communication overhead by up to 95%. By separating compression features in frequency and spatial domain, we propose a new defense mechanism that can reduce the transmission overhead of the attacked input by about 9% compared to unperturbed data, with only about 2% accuracy loss. Importantly, the proposed defense mechanism is a standalone approach which can be applied in conjunction with approaches such as adversarial training to further improve robustness. Code is available at https://github.com/Restuccia-Group/EntropyR.
@@ 2 Entropy Coding in Distributed DNNs
Recent work [42] has incorporated entropy coding into distributed DNNs to increase communication efficiency between mobile and edge devices. The proposed method, referred to as the entropic student, aims at achieving a balance between the coding rate and task-specific performance metrics (e.g., cross entropy in classification tasks). One distinction between the entropic student and other entropy-coding-based feature compression work [55, 59] lies in the significantly smaller encoder size of the former due to constraints on computational resources in mobile devices. As a result, the entropic student splits the DNN at an early layer and use knowledge distillation [26] for training the head to preserve accuracy. However, existing work does not address the issue of resilience of entropy models, which is the key goal of this paper. A relevant work [4] investigated the robustness of density estimation aiming to maximize $P_{Z}(z)$ while we are interested in perturbations that minimize $P_{Z}(z)$.
@@ 3 Modeling Interference to Entropy Coding
$\min_{\delta}$ $\log_{2}P_{Z}(H(x+\delta))$ (7)
$s.t.$ $||\delta||_{p}\leq\epsilon$
@@ 7 Conclusion
Acknowledgements
This work is funded in part by the Air Force Office of Scientific Research under contract number FA9550-23-1-0261, by the Office of Naval Research under contract number N00014-23-1-2221, as well as by National Science Foundation grants CNS-2134973 and CNS-2312875. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the U.S. Government.